# Outlier-robust Kalman Filtering through Generalised Bayes

### v1

@@ 1 Introduction
Probabilistic state-space models (SSMs) are widely used to address problems in time-series forecasting, online learning, tracking problems, and signal processing. A key challenge in SSMs is to perform online (sequential) posterior inference, also known as Bayesian filtering. If the model is linear and Gaussian, then the Kalman filter (KF) algorithm is the optimal filter in terms of minimal mean squared error (Morris, 1976). If the model is not linear or not Gaussian, approximate inference methods must be used. Although there are a number of techniques available, in this paper we focus on the extended Kalman filter (EKF) (see e.g., Sarkka & Svensson, 2023) and the ensemble Kalman filter (EnKF) (see e.g., Roth et al., 2017a), because they both admit closed-form Bayesian updates to Gaussian posterior approximations and scale to high dimensions.
@@
The KF and KF-like methods typically use a Gaussian observation model for computational convenience. However, a weakness of this assumption is that extreme observations, such as outliers, are not well modelled by Gaussians, leading to model misspecification (Grubbs, 1969; Hampel et al., 2011). Moreover, in practice, the true dynamics are often different from the one assumed by the filter. This could be because the measurements are corrupted, or because the user does not have access to the measurement function. For example, in tracking problems, sensor measurement errors record the wrong position of an object; similarly, in economics and finance, there are instances of data providers sending time series with erroneous values (see e.g. Liu, 2020, for a survey).
There is a large literature that studies the filtering problem in the presence of outliers and misspecified measurement models [^1] Alternatively, a measurement model can be referred to as an observation model.. Some approaches extend the state space by introducing hierarchical priors on the measurement model. Since closed-form estimation of the state is not generally feasible, a factorised variational Bayes (VB) approach can be employed to obtain a fixed-point solution (see e.g., Ting et al., 2007; Agamennoni et al., 2012; Huang et al., 2016; Nurminen et al., 2015; Wang et al., 2018; Piché et al., 2012). Alternatively, huberised (E)KF-type methods are optimisation-based approaches to filtering, which minimise the Huber loss; see Boncelet & Dickinson (1983); Karlgaard (2015); Das (2023) and the references therein. Recently, Boustati et al. (2020) proposed a particle filter (PF) approach that uses generalised Bayes (GB) to handle model misspecification from outliers; this builds on prior work showing that GB updates are theoretically sound and robust to outliers and model misspecification (see, e.g., Bissiri et al., 2016; Jewson et al., 2018; Knoblauch et al., 2022; Fong et al., 2021; Jewson & Rossell, 2022; Husain & Knoblauch, 2022; Matsubara et al., 2023, 2022). However, these previous approaches can be slow, due to the need to perform iterative optimisation per step (for VB) (Knoblauch et al., 2018), or the need to use a large number of samples (for PF) (Boustati et al., 2020). See Appendix A for an overview of a number of these methods.
@@ 3.1 Linear weighted observation likelihood filter
The computational complexity of WoLF under linear dynamics matches that of the KF, i.e., $O(m^{3})$.333 This can be reduced to $O(\min\{dm^{2},d^{2}m\})$ when $d<m$ by rewriting the precision update in terms of variance using the Woodbury identity. Alternative robust filtering algorithms require multiple iterations per measurement to achieve robustness and stability, making them significantly slower; see Table 1 for the computational complexity for the methods we consider, and Table 2 and Figure 5 for empirical comparisons.
@@ 3.3 The choice of weighting function
For ${\bm{y}}_{t}\in{}^{d}$ with $d=m\gg 1$ and diagonal measurement covariance $\mathbf{R}_{t}=\text{diag}\left(r_{t,1},\ldots,r_{t,d}\right)$, the WoLF-TMD function can be modified to weight individual observations so that $W:{}^{2d}\to{}^{d}$. See Section 4.3 for an example and Appendix D.2 for a discussion.
The proposed weighting functions — the IMQ, the MD, and the TMD — are defined such that $W:{}^{2d}\to[0,1]$ and therefore can only down-weight observations. This means that our updates are always conservative, i.e., our posteriors will be wider in the presence of outliers (see Figure 1 for an example).
@@ Results
Figure 3 shows a sample of each variant along with the filtered state for each method. For the Student variant (left panel), the WoLF-IMQ and the WoLF-TMD estimate the true state more closely than the competing methods. Both the KF-IW and the KF-B look comparable to the KF, which are not robust to outliers. For the mixture variant (right panel),444 The top left figure in the right panel is cropped, see Figure 13 for the uncropped version. we see that the WoLF-IMQ, the WoLF-TMD, and the KF-B filter the true state correctly. In contrast, the KF-IW and the KF are not robust to outliers.555 KF-B removes outliers that bear no information according to some criterion, but in the Student-t case, it fails. KF-IW, on the other hand, estimates a measurement covariance rather than the dispersion of such a measurement covariance. In this sense, it is misspecified in both cases.
@@ 4.2 Robust EKF for online MLP regression (UCI)
In this section, we benchmark the methods using a corrupted version of the tabular UCI regression datasets. [^6] The dataset is available at https://github.com/yaringal/DropoutUncertaintyExps. Similar to other papers that deal with non-linear online learning (see, e.g. Chang et al., 2023), we consider a single-hidden-layer multi-layered perceptron (MLP) with twenty hidden units and a real-valued output unit. In this experiment, the state dimension (number of parameters in the MLP) is $m=(n_{\text{in}}\times 20+20)+(20\times 1+1)$, where $n_{\text{in}}$ is the dimension of the feature ${\bm{x}}_{t}$.777 See Table 3 for the values that $n_{\text{in}}$ takes for each dataset. One of the main advantages of using a Bayesian filtering method for fitting neural networks (compared to using OGD) is the ability to handle non-stationary distributions (see e.g. Chang et al., 2023). Below, we take a static state (see e.g. Lambert et al., 2021), so that the prior predictive mean is $\bm{\mu}_{t|t-1}=\bm{\mu}_{t-1}$. In Appendix E.3 we study online learning with non-stationary environments.
@@ 5 Conclusion
6 Impact statement
@@ 7 Acknowledgements
Appendix A Background on existing robust filters
@@
To the best of our knowledge, the first work that proposes robust filtering in the context of robust statistics is Calvet et al. (2015). Their robust filter follows Masreliez & Martin (1977) and is based on a “Huberisation” of the derivative of the log-measurement density (score function), which then they integrate. See Schick & Mitter (1994) for a comprehensive review of classical robust-KF methods for linear SSMs.
Recently proposed provably-robust methods include the work by Boustati et al. (2020) and Cao et al. (2022).
@@ D.2 Dimension-specific weighting
In this section, we provide details about the implementation of a vector-valued weighting function $W_{t}:{}^{2d}\to{}^{d}$ introduced in Section 3.3. See Section 4.3 for an evaluation of this method. We consider ${\bm{y}}_{t}\in{}^{d}$ with $d>1$.

### v2

@@ 1 Introduction
Probabilistic state-space models (SSMs) are widely used to address problems in time-series forecasting, online learning, tracking problems, and signal processing. A key challenge in SSMs is to perform online (sequential) posterior inference, also known as Bayesian filtering. If the model is linear and the and the dynamics are known, then the Kalman filter (KF) algorithm is the optimal filter in terms of minimal mean squared error (Morris, 1976). If the model is not linear, approximate inference methods must be used. Although there are a number of techniques available, in this paper we focus on the extended Kalman filter (EKF) (see e.g., Sarkka & Svensson, 2023) and the ensemble Kalman filter (EnKF) (see e.g., Roth et al., 2017a), because they both admit closed-form Bayesian updates to Gaussian posterior approximations and scale to high dimensions.
@@
In a Bayesian setting, the KF and KF-like methods typically use a Gaussian measurement model for computational convenience. However, a weakness of this assumption is that extreme observations, such as outliers, are not well modelled by Gaussians, leading to model misspecification (Grubbs, 1969; Hampel et al., 2011). Moreover, in practice, the true dynamics are often different from the one assumed by the filter. This could be because the measurements are corrupted, or because the user does not have access to the measurement function. For example, in tracking problems, sensor measurement errors record the wrong position of an object; similarly, in economics and finance, there are instances of data providers sending time series with erroneous values (see e.g. Liu, 2020, for a survey).
There is a large literature that studies the filtering problem in the presence of outliers and misspecified measurement models [^1] Alternatively, a measurement model can be referred to as an observation model.. Some approaches extend the state space by introducing hierarchical priors on the measurement model. Since closed-form estimation of the state is not generally feasible, a factorised variational Bayes (VB) approach can be employed to obtain a fixed-point solution (see e.g., Ting et al., 2007; Agamennoni et al., 2012; Huang et al., 2016; Nurminen et al., 2015; Wang et al., 2018; Piché et al., 2012). Alternatively, Huberised (E)KF-type methods are optimisation-based approaches to filtering, which minimise the Huber loss; see Boncelet & Dickinson (1983); Karlgaard (2015); Das (2023) and the references therein. Recently, Boustati et al. (2020) proposed a particle filter (PF) approach that uses generalised Bayes (GB) to handle model misspecification from outliers; this builds on work showing that GB updates are theoretically sound and robust to outliers and model misspecification (see, e.g., Bissiri et al., 2016; Jewson et al., 2018; Knoblauch et al., 2022; Fong et al., 2021; Jewson & Rossell, 2022; Husain & Knoblauch, 2022; Matsubara et al., 2023, 2022). However, these approaches can be slow, due to the need to perform iterative optimisation per step (for VB) (Knoblauch et al., 2018), or the need to use a large number of samples (for PF) (Boustati et al., 2020). See Appendix A for an overview of a number of these methods.
@@ 3.1 Linear weighted observation likelihood filter
The computational complexity of WoLF under linear dynamics matches that of the KF, i.e., $O(m^{3})$. Alternative robust filtering algorithms require multiple iterations per measurement to achieve robustness and stability, making them significantly slower; see Table 1 for the computational complexity for the methods we consider, and Table 2 and Figure 5 for empirical comparisons.
@@ 3.3 The choice of weighting function
For ${\bm{y}}_{t}\in{}^{d}$ with $d=m\gg 1$ and diagonal measurement covariance $\mathbf{R}_{t}=\text{diag}\left(r_{t,1},\ldots,r_{t,d}\right)$, the WoLF-TMD function can be modified to weight individual observations so that $W:{}^{d}\times{}^{d}\to{}^{d}$. See Section 4.3 for an example and Appendix D.2 for a discussion.
The proposed weighting functions — the IMQ, the MD, and the TMD — are defined such that $W:{}^{d}\times{}^{d}\to[0,1]$ and therefore can only down-weight observations. This means that our updates are always conservative, i.e., our posteriors will be wider in the presence of outliers (see Figure 1 for an example).
@@ Results
Figure 3 shows a sample of each variant along with the filtered state for each method. For the Student variant (left panel), the WoLF-IMQ and the WoLF-TMD estimate the true state more closely than the competing methods. Both the KF-IW and the KF-B look comparable to the KF, which are not robust to outliers. For the mixture variant (right panel),333 The top left figure in the right panel is cropped, see Figure 13 for the uncropped version. we see that the WoLF-IMQ, the WoLF-TMD, and the KF-B filter the true state correctly. In contrast, the KF-IW and the KF are not robust to outliers.444 KF-B removes outliers that bear no information according to some criterion, but in the Student-t case, it fails. KF-IW, on the other hand, estimates a measurement covariance rather than the dispersion of such a measurement covariance. In this sense, it is misspecified in both cases.
@@ 4.2 Robust EKF for online MLP regression (UCI)
In this section, we benchmark the methods using a corrupted version of the tabular UCI regression datasets. [^5] The dataset is available at https://github.com/yaringal/DropoutUncertaintyExps. Similar to other papers that deal with non-linear online learning (see, e.g. Chang et al., 2023), we consider a single-hidden-layer multi-layered perceptron (MLP) with twenty hidden units and a real-valued output unit. In this experiment, the state dimension (number of parameters in the MLP) is $m=(n_{\text{in}}\times 20+20)+(20\times 1+1)$, where $n_{\text{in}}$ is the dimension of the feature ${\bm{x}}_{t}$.666 See Table 3 for the values that $n_{\text{in}}$ takes for each dataset. One of the main advantages of using a Bayesian filtering method for fitting neural networks (compared to using OGD) is the ability to handle non-stationary distributions (see e.g. Chang et al., 2023). Below, we take a static state (see e.g. Lambert et al., 2021), so that the prior predictive mean is $\bm{\mu}_{t|t-1}=\bm{\mu}_{t-1}$. In Appendix E.3 we study online learning with non-stationary environments.
@@ 5 Conclusion
Acknowledgements
JK and FXB were supported by the EPSRC grant [EP/Y011805/1].
Impact statement
@@ Acknowledgements
Appendix A Background on existing robust filters and related methods
@@
To the best of our knowledge, the first work that proposes robust filtering in the context of robust statistics is Calvet et al. (2015). Their robust filter follows Masreliez & Martin (1977) and is based on a “Huberisation” of the derivative of the log-measurement density (score function), which then they integrate. See Schick & Mitter (1994) for a comprehensive review of classical robust-KF methods for linear SSMs. Recently proposed provably-robust methods include the work by Boustati et al. (2020) and Cao et al. (2022).
Finally, we discuss similarities and differences with the Bayesian learning rule (BLR) line of work by Khan & Rue (2023). BLR also allows the replacing of the likelihood with a loss function, but that loss depends only on the parameter and the observation, i.e., BLR uses $\ell_{t}({\bm{\theta}})=L({\bm{y}}_{t},f_{t}({\bm{\theta}}))$, with $L:{}^{d}\times{}^{d}$ a loss function. By contrast, our loss depends on the weight $W_{t}$, which in turn depends on the belief state, via $\hat{{\bm{y}}}_{t}=f_{t}(\bm{\mu}_{t|t-1})$, which in turn is a function of all the past data, i.e., our loss has the form $\ell_{t}({\bm{\theta}})=L({\bm{y}}_{1:t},f_{t}({\bm{\theta}}))$. In addition, WoLF is designed for the online inference setting, whereas BLR, at least in its standard form, is designed for offline inference. Finally, note that the exponential family extension of BLR Lin et al. (2019) has also been proposed as a way to get robustness by using scale mixture posteriors. However, this induces robustness in parameter space, whereas we focus on robustness in observation space. In particular, the weighting term for the BLR mixture model depends on the distance in parameter space, ${\bm{\theta}}_{t}-\bm{\mu}_{t}$, whereas ours depends on the difference in observation space, ${\bm{y}}_{t}-\hat{{\bm{y}}_{t}}$.
@@ D.2 Dimension-specific weighting
In this section, we provide details about the implementation of a vector-valued weighting function $W_{t}:{}^{d}\times{}^{d}\to{}^{d}$ introduced in Section 3.3. See Section 4.3 for an evaluation of this method. We consider ${\bm{y}}_{t}\in{}^{d}$ with $d>1$.